Suppose Lauren, Leslie and Lydia all purchase small white board markers for their rooms at a certain price each. Lauren's willingness to pay was R350, Leslie's willingness to pay was R250, and Lydia's willingness to pay was R300. Total consumer surplus for these three was R450. What is the price of each white board marker? [1] R150. [2] R200. [3] R950. [4] R1350. 


1
Expert's answer
2021-04-11T20:03:06-0400

If P is market price, then total consumer surplus is:

CS = (350 - P) + (300 - P) + (250 - P) = 450.

900 - 3P = 450,

P = R150.

So, the correct answer is [1] R150.